OpenStudy (anonymous):

Solving integrals. Please explain how you got the answer. Thank you! 1. ∫e^(-2x)*(3+4e^(-2x))^6 dx 2. ∫ x/(9+16x^2) dx 3. ∫ (3x+1)/sqrt(5x^2+3) dx 4. ∫ x/sqrt(5+4x-x^2) dx

OpenStudy (anonymous):

\[I=\int\limits e ^{-2x}*\left\{ 3+4e ^{-2x} \right\}^{6}dx\] \[put~ 3+4e ^{-2x}=y,e ^{-2x}=\frac{ y-3 }{4 },e ^{-2x}\left( -2 \right)dx=\frac{ 1 }{ 4 }dy\] \[e ^{-2x}dx=-\frac{ 1 }{ 8 }dy\] \[I=-\frac{ 1 }{ 8 } \int\limits y^6 dy\] complete it.

OpenStudy (anonymous):

4. ∫ x/sqrt(5+4x-x^2) dx \[(1)~~\int e^{-2x}\left(3+4e^{-2x}\right)^6~dx\] Substitute \(u=3+4e^{-2x}\), then \(du=-8e^{-2x}~dx\), or \(-\dfrac{1}{8}du=e^{-2x}~dx\): \[-\frac{1}{8}\int u^6~du=\cdots\] - - - - - - - - - - - - \[(2)~~\int\frac{x}{9+16x^2}~dx\] Substitute \(t=9+16x^2\), then \(dt=32x~dx\), or \(\dfrac{1}{32}dt=x~dx\): \[\frac{1}{32}\int\frac{dt}{t}~dt=\cdots\] - - - - - - - - - - - - \[(3)~~\int\frac{3x+1}{\sqrt{5x^2+3}}~dx=3\int\frac{x}{\sqrt{5x^2+3}}~dx+\int\frac{dx}{\sqrt{5x^2+3}}\] For the first integral, substitute \(p=5x^2+3\), then \(dp=10x~dx\), or \(\dfrac{1}{10}dp=x~dx\). For the second integral, substitute \(x=\sqrt{\dfrac{3}{5}}\tan r\), then \(dx=\sqrt{\dfrac{3}{5}}\sec^2r~dr\). \[\frac{3}{10}\int\frac{dp}{\sqrt{p}}+\int\frac{\sqrt{\frac{3}{5}}\sec^2r}{\sqrt{5\left(\sqrt{\frac{3}{5}}\tan r\right)^2+3}}~dr\\ \frac{3}{10}\int p^{-1/2}~dp +\sqrt{\frac{3}{5}}\int\frac{\sec^2r}{\sqrt{5\left(\frac{3}{5}\tan^2 r\right)+3}}~dr\\ \frac{3}{10}\int p^{-1/2}~dp +\sqrt{\frac{3}{5}}\int\frac{\sec^2r}{\sqrt3\sqrt{\tan^2 r+1}}~dr\\ \frac{3}{10}\int p^{-1/2}~dp +\frac{1}{\sqrt{5}}\int\frac{\sec^2r}{\sqrt{\sec^2 r}}~dr\\ \frac{3}{10}\int p^{-1/2}~dp +\frac{1}{\sqrt{5}}\int \sec r~dr=\cdots\] - - - - - - - - - - - - \[(4)~~\int\frac{x}{\sqrt{5+4x-x^2}}~dx\] Complete the square in the denominator: \[\begin{align*}5+4x-x^2&=-\left(x^2-4x-5\right)\\ &=-\left(x^2-4x+4-9\right)\\ &=-\left((x-2)^2-9\right)\\ &=9-(x-2)^2\end{align*}\] \[\int\frac{x}{\sqrt{9-(x-2)^2}}~dx\] Substitute \(w=x-2\), then \(dw=dx\): \[\int\frac{w+2}{\sqrt{9-w^2}}~dw=\int\frac{w}{\sqrt{9-w^2}}~dw+2\int\frac{dw}{\sqrt{9-w^2}}\] For the first integral, substitute \(q=9-w^2\), then \(dq=-2w~dw\), or \(-\dfrac{1}{2}dq=w~dw\). For the second integral, substitute \(w=3\sin z\), then \(dw=3\cos z~dz\): \[-\frac{1}{2}\int\frac{dq}{\sqrt{q}}+2\int\frac{3\cos z}{\sqrt{9-(3\sin z)^2}}~dz\\ -\frac{1}{2}\int q^{-1/2}~dq+\frac{6}{\sqrt9}\int\frac{\cos z}{\sqrt{1-\sin^2z}}~dz\\ -\frac{1}{2}\int q^{-1/2}~dq+\frac{6}{\sqrt9}\int\frac{\cos z}{\sqrt{\cos^2z}}~dz\\ -\frac{1}{2}\int q^{-1/2}~dq+\frac{6}{\sqrt9}\int dz=\cdots\]